# RING-NeRF: A Versatile Architecture based on Residual Implicit Neural Grids

### v1

@@ 4.2 Novel View Synthesis and Anti-Aliasing
Nerfacto 27.09 0.779 0.181 0.18 h
Zip-NeRF (repr.) 28.06 0.808 0.154 0.89 h
Our Model 28.11 0.802 0.155 0.18 h
Table 1: Novel View synthesis performances for the Mono-Scale setup on the 360 dataset.
PSNR ↑ SSIM ↑ LPIPS ↓
Image Resolution Full Res. 1/2 Res. 1/4 Res. 1/8 Res. Full Res. 1/2 Res. 1/4 Res. 1/8 Res. Full Res. 1/2 Res. 1/4 Res. 1/8 Res. Time ↓
Mip-NeRF 360 27.51 29.19 30.45 30.86 0.779 0.864 0.912 0.931 0.254 0.136 0.077 0.058 21.86h
Nerfacto 25.98 27.31 27.16 25.30 0.668 0.792 0.805 0.743 0.367 0.176 0.139 0.199 0.20h
Zip-NeRF (repr.) 27.54 29.34 30.36 30.90 0.781 0.858 0.889 0.900 0.191 0.099 0.064 0.047 0.90h
Our Model 27.67 29.50 30.48 30.86 0.780 0.858 0.891 0.903 0.196 0.092 0.061 0.049 0.20h
Table 2: Novel View Synthesis performances for the Multi-Scale setup on the 360 Dataset.
@@
Results. Regarding quality for the mono-scale setup, as referred in table 1, we notice that our architecture provides the best performances. RING-NeRF is on par with Zip-NERF, slightly better than Mip-NeRF-360, and even better than Nerfacto. Regarding quality for the multi-scale setup, as referred in table 2 and illustrated in figure 1, we observe that, on one side, all the algorithms that consider the distance of observation perform similarly in terms of quality [^4] Please note that in the paper Zip-NeRF [3], slightly better performances are referred in mono scale (PSNR of 28.54) and multi-scale (PSNR of 28.25/30/31.57/32.52) setups. However, we were not able to reproduce those results, probably because of a difference in the evaluation pipeline and implementation..
On the other side, Nerfacto, which does not take scale and distances into account when processing a sample, performs poorly on the dataset. Its performances are particularly low at the coarsest resolution, with over-contrasted images and even aliasing artifacts such as the ones illustrated in figure 3.
Regarding the reconstruction and rendering processing time, Nerfacto and our architecture perform similarly both on mono-scale and multi-scale setups. Zip-NeRF is second with a processing time approximately 5 times slower, and MIP-NeRF-360 is extremely slow since based on vanilla NeRF architecture.
In conclusion, our solution provides the best quality-speed trade-off since it is both on par with the best quality method and the fastest method, in mono-scale as well as in multi-scale setups. Zip-NeRF is equivalent in terms of quality in the multi-scale setup but with a processing time 5 times slower due to its use of super-sampling.
4.3 Few Viewpoints Supervision
PSNR ↑ SSIM ↑ LPIPS ↓ Time ↓
Mip-NeRF 8.68 / 16.54 / 23.58 0.571 / 0.741 / 0.879 0.353 / 0.198 / 0.092 2.56h
FreeNeRF 19.92 / 23.25 / 25.38 0.787 / 0.844 / 0.888 0.182 / 0.137 / 0.096 2.56h
Nerfacto 9.35 / 9.75 / 9.78 0.567 / 0.604 / 0.647 0.385 / 0.331 / 0.326 0.15h
Nerfacto+ 13.61 / 16.61 / 19.33 0.639 / 0.699 / 0.759 0.276 / 0.218 / 0.151 0.15h
Our Model 16.18 / 20.47 / 23.19 0.713 / 0.808 / 0.847 0.200 / 0.127 / 0.085 0.15h
w/ discrete CtF 15.79 / 20.16 / 22.93 0.706 / 0.785 / 0.847 0.201 / 0.127 / 0.085 0.15h
Table 3: Performances of reconstruction from few viewpoints (respectively 3, 6 and 9, separated by ”/”) on the DTU dataset. The reported metrics are computed based on the mask of the object.
Algorithms. We compare our architecture against several baselines: Mip-NeRF [1] and FreeNerf [19], one state-of-the-art method for this task, for vanilla architectures and Nerfacto for grid-based architectures. Since, to the best of our knowledge, there is no grid-based solution for this task, we developed the Nerfacto+ architecture, which corresponds to a Nerfacto architecture coupled with a coarse-to-fine training based on a progressive activation of the hash grids (the rest of the decoder’s input being filled with zeros). For Nerfacto+ and RING-NeRF, we also add the FreeNeRF’s loss that penalizes the density first $M=10$ samples of each ray to reduce as most as possible artifacts in front of the cameras. In order to demonstrate the intrinsic stability of RING-NeRF, we also evaluate the model using discrete coarse-to-fine (using a fixed LOD increment of 1 rather than the proposed continuous increase) as an ablative experiment.
Protocol. We follow FreeNeRF’s evaluation pipeline, including the number of supervision viewpoints (3 to 9), the choice of these views among the dataset and the evaluations of metrics using masks of the object and we also deactivate the space compression. Since we are following the FreeNeRF evaluation protocol, the results of FreeNeRF and Mip-NeRF were taken out of FreeNeRF’s article.
Results. Evaluation results are referred in table 3. We first notice an important difference between vanilla and grid-based baselines. While Mip-NeRF faces troubles in reconstructing the scene with 3 views, the method seems to find coherency when adding a few more images. However, Nerfacto struggles much more to form a consistent 3D scene even when using 9 images (see figure 1). While the grid-based baseline is way faster to train, its design seems to imply more instability when facing a small number of images.
This does not mean however that few viewpoints are incompatible with grid-based methods. Using progressive training coupled with a very simple and generalizable regularization, both Nerfacto+ and our architecture succeed in creating 3D geometry. Nonetheless, RING-NeRF considerably outperforms Nerfacto+, with a PSNR difference varying from 3 to 4, demonstrating the stability increase brought by our architecture, and also outperforms Mip-NeRF for the configuration with 3 and 6 supervision images. The discrete ablative version of our model places in-between Nerfacto+ and our complete RING-NeRF, showcasing both the intrinsic interest of our architecture and the relevance of the continuous coarse-to-fine mechanism. FreeNerf remains the best performer of all in terms of quality, but with a reconstruction time that is extremely slower than RING-NeRF, the former achieving its reconstruction in 2.56 hours while the latter only requires less than 10 minutes. RING-NeRF thus offers a better quality-speed trade-off (see figure 1). Beyond the metrics, we also notice experimentally more instability on the other grid-based architecture, resulting in more frequent training divergences when slightly tweaking the hyperparameters on Nerfacto+.
4.4 SDF Reconstruction without Initialization
SDF reconstruction is known to be a more unstable process than density NeRF reconstruction, requiring a scene-specific initialization to converge. This scene-specific initialization becomes an issue in complex environments and incremental setups, where several types of scenes can co-exist and are not necessarily known beforehand. Therefore, in this experiment, we evaluate the ability of RING-NeRF and other SotA architectures to achieve SDF reconstruction without scene-specific initialization.
Dataset. The evaluation is achieved on a subset of 7 scenes of the Replica[15] synthetic indoor dataset.
Algorithms. We compare our architecture to two SDF methods, all of them implemented in the same SDFStudio [20] branch of the Nerfstudio framework for fairer comparisons: NeuS-facto, an adaptation of NeuS for grid-based methods and an implementation of NeuralAngelo. We train each model for both 20k epochs and 100k epochs with appropriate schedulers.
PSNR ↑ SSIM ↑ LPIPS ↓ Training Time ↓
NeuS-facto (20k ep.) 19.45 0.721 0.722 0.27 h
NeuS-facto (100k ep.) 24.62 0.778 0.347 1.4h
NeuralAngelo (20k ep.) (repr.) 25.09 0.830 0.206 0.67 h
NeuralAngelo (100k ep.) (repr.) 30.79 0.916 0.0761 3.40 h
Our Model (20k ep.) 34.37 0.934 0.0546 0.25 h
Our Model (100k ep.) 37.18 0.969 0.0194 1.28 h
Table 4: SDF reconstruction performances when foregoing the scene-specific SDF Initialization on the Replica Dataset.
Results. Evaluation results are referred in table 4. First of all, the NeuS-facto method faces low PSNR in mean, due to catastrophic failure in most tested scenes (see figure 1) since the model quickly tends to re-draw the 2D images in front of the camera. Regarding NeuralAngelo, its relatively high PSNR demonstrates its ability to synthesize satisfying RGB. However, as illustrated in figure 1, the underlying geometry of the scene is poorly reconstructed, without any fine details. Finally, our method RING-NeRF is by far the best performer, with much higher PSNR and a better geometry including fine details (see figure 1), although a bit noisy.
The simple architecture of RING-NeRF also permits faster epochs, thus faster training.
4.5 Generalization and LOD extensibility
PSNR ↑ SSIM ↑ LPIPS ↓
Free (3LvLs) 27.28 0.814 0.308
Frozen (3LvLs) 27.03 0.789 0.332
Free (4LvLs) 28.58 0.849 0.243
Frozen (+1LvL) 28.43 0.849 0.248
Free (5LvLs) 29.76 0.884 0.153
Frozen (+2LvLs) 29.66 0.875 0.175
Table 5: Reconstruction performances on one DTU scene with both the usual setup (free decoder) and different increases of resolution with a frozen decoder.
This experiment aims to demonstrate both the generalization ability of RING-NeRF when using a frozen pre-trained decoder, but also its unique ability to increase dynamically the level of details of the scene representation.
Protocol. We pre-train the decoder of our model through the joint reconstruction of several DTU scenes with separate grids but one shared decoder. During the pre-training, only 3 levels of grids were used, with respective resolutions of $16^{3}$, $32^{3}$ and $64^{3}$.
A DTU scene that was not used during the pre-training stage is then reconstructed with the usual pipeline, except that only the grids parameters are optimized with the locked pre-trained decoder. A first reconstruction is achieved with 3 levels of grids. The 3 grids are then frozen and two other reconstructions are launched: the first one adding one more grid level( resolution $128^{3}$) and another one adding two grid levels (resolution $128^{3}$ and $256^{3}$), those grids being initialized with zero. The performances are compared to a classic RING-NeRF reconstruction, with the decoder and dense grids being initialized and refined as usual.
Results. Performances are reported in table 5. First, we notice that using a pre-trained decoder has little impact on the performances, demonstrating the ability of a pre-trained RING-NeRF to generalize on new scenes. Second, we observe that more levels of grids improve the performances, even if the additional grid resolutions were not in the pre-training, and that the 3 first levels of grids are frozen. It demonstrates the ability of RING-NeRF to increase dynamically its level of detail, a feat that most grid-based architectures such as I-NGP cannot perform, even when pre-trained as the size of the decoder’s input is fixed.
In our opinion, these results demonstrate the benefits of having a decoder latent space invariant to the position and LOD of the samples. It also shows that RING-NeRF can achieve progressive reconstruction, a coarse LOD being available for consultation while finer LOD being optimized. This property is very useful for both incremental and adaptive reconstruction processes.
5 Conclusion and perspectives
In this work, we introduced RING-NeRF, a versatile NeRF pipeline based on Residual Implicit Neural Grids that provides control on the level of detail of the mapping function between the scene and the latent space. Including a distance-aware forward mapping and a coarse-to-fine reconstruction process, our pipeline demonstrated experimentally on-par performances with specifically designed state-of-the-art solutions for anti-aliasing or reconstruction from few supervised viewpoints without slowing down the reconstruction process nor limiting it to object-centric scenes and even outperforms them in terms of robustness to scene-specific initialization for SDF reconstruction.
Future work will study the impact of RING-NeRF on other challenging use cases, such as facing inaccurate camera poses and SLAM [23]. We will also take advantage of the extensibility property of our architecture to develop memory-efficient sparse grid-based Neural Fields.
Acknowledgements
This publication was made possible by the use of the CEA List FactoryIA supercomputer, financially supported by the Ile-de-France Regional Council
Supplementary Material
Figure 4: Different LODs Outputs of the model when only trained on the last LOD. We observe that, even without supervising intermediate LODs during the training, a notion of LOD is captured in the scene reconstruction. We also observe visually continuous LOD since, as expected, the LOD 3.5 outputs a 3D representation in between LOD 3 and 4 in term of details.
Figure 5: Comparison of LOD when supervising (a) every level of detail or (b) solely the last level of detail
6 Architecture
6.1 Architecture details
As mentioned in the article, the final feature after multi-resolution combination is passed into a normalization layer. Novel View Synthesis experiments showed slightly better results when using standard, non-learnable normalization over the feature. This is followed by the learnable frequency filter, for which we chose a sinus filter : $y=sin(Wx)$ with $W$ a linear matrix (without bias). A linear layer transforms the filter output into both density (or SDF) and one color feature. This feature, concatenated with the direction of observation encoded into Spherical Harmonics, is fed to a MLP of 3 layers that predicts the radiance.
6.2 Differences with other architectures
A first noticeable difference is related to the notion of LOD that is induced by the architecture itself, independently to any supervision. This property is justified through the interpretation of the grid as a mapping function from the scene space to the decoder latent space (see section 3.2), but is also validated experimentally (see section 6.3 and figure 4). Except TriMip-RF [6] and LoD-NeuS [24] which possess such property but for tri-plan, other architectures need an explicit supervision to capture the notion of LOD (see ResidualMFN  [^5] S. Shekarforoush, D. Lindell, D. Fleet and M. Brubaker, Residual Multiplicative Filter Networks for Multiscale Reconstruction, in: Advances in Neural Information Processing Systems, 2022. for instance). This property is probably one of the main reason of the convergence robustness of RING-NeRF, as observed through few-view reconstruction and SDF reconstruction experiments. It also simplifies the setup of a coarse-to-fine reconstruction process and a distance-aware forward mapping without scarifying speed since no convolution [24] nor super-sampling [3] is implied.
A second difference is the invariance of the decoder latent space towards position (unlike NGLOD [16], Tri MipRF [6], etc, that concatenate Positional Encoding with the feature extracted from the grid) but also level of detail. LOD is exclusively encoded in the grid, unlike NGLOD [16] or MFLOD [4] that use a per LOD decoder, or Zip-NeRF [3] and Mip-NeRF [1] which use down-weighting. This property facilitates the pre-training of the decoder, since the learnt latent space is independent of the scale and position of the scene elements, and then favor the generalization of a pre-trained decoder to new scene.
Finally, the last specificity is related to the combination of the independence of the decoder latent space with respect to the size of the grid hierarchy, and the top-down representation of the LOD. As we confirmed experimentally (see section 4.5 of the article), this combination makes the maximal level of detail unbounded since the size of the grid pyramid can grow up dynamically (similarly to [11]) while keeping the access to the different LODs available (unlike [11]).
6.3 Architecture-induced LOD
As discussed in section 6.2, an interesting intrinsic property of the RING-NeRF architecture is that it naturally produces Level Of Detail of the reconstruction, even when we only use the full resolution images to supervise the full resolution LOD and not a pyramid of images supervising corresponding intermediate LOD like usual LOD NeRF architectures. This property is illustrated on figure 4 that shows, for several scenes of the DTU dataset, different LOD outputs obtained from a same reconstruction pipeline that is supervised only in the last level of detail. While not providing as good results as those obtained (see figure 5) by supervising all the LOD, we observe however a correlation between the LOD used to compute the image and the level of details in the images. It confirms experimentally that our architecture induces by itself a notion of LOD in the 3D representation.
7 New view synthesis - Section 4.2
7.1 Configuration
Our Novel View Synthesis and Anti-Aliasing experiments on the 360 dataset are done on one unique configuration for our baselines Nerfacto, ZipNeRF reproduction and for our model RING-NeRF, both for the mono-scale and multi-scale setups. We mostly used the configuration provided in the article ZipNeRF with few differences that we will underline in the following. Each of these three baselines were trained for 100k iterations with a batch size of $2^{14}$ rays. We use the RAdam Optimizer with an initial learning rate of $21e-$ 2, and a cosine-decayed scheduler over the whole training to $31e-$ 3. Both Nerfacto and RING-NeRF discard the proposed warm-up for an increase in stability.
Regarding the model itself, we also use a hash grid pyramid of 10 grids with a growth factor of 2, with grid resolutions ranging from 16 to 8192. Contrary to the original Zip-NeRF, we use 8 features per level rather than 4, but with a similar hashmap size of $128^{3}=2^{21}$. We also use proposal samplers introduced by mip-NeRF 360, and once again follow Zip-NeRF configuration : 2 rounds of sampling via 2 different grid-based proposal samplers (hashgrid pyramid + one linear layer), and one last forward process into the true model to generate the pixel values. Our 2 samplers have respectively 512 and 2048 max resolution and both use features of size 1 as the sampling process only needs density information. For simplicity of implementation and comparison, we actually use the usual I-NGP-based concatenation model for the proposal samplers.
Following Nerfacto’s architecture, we also used scene contraction, as well as the distortion loss in these experiments, as implemented in NerfStudio for both Nerfacto and RING-NeRF, and their adapted implementations versions for ZipNeRF. We also did not use appearance embedding as we also noticed a decrease in metrics when using it, as observed in Mip-NeRF 360 and ZipNeRF. Finally, for both Nerfacto and RING-NeRF, we did not use any additional mechanism such as the ones introduced in ZipNeRF (no novel interlevel loss, nor scale featurization, nor weight decay loss, nor Affine Generative Latent Optimization as appearance embedding, …).
Figure 6: Example of effect of the distance-aware mechanism on the reconstruction of distant areas.
7.2 Result Analysis and Ablative Experiments
Mono-Scale Setups. In order to further analyze the separate impact of our architecture and of the associated mechanisms (distance-aware forward mapping and continuous coarse-to-fine), we also run ablative mono-scale experiments. Our model is thus evaluated on 4 configurations with varying setups and results can be found in table 6.
First of all, we notice a gap between the results of Nerfacto (27.09, 0.779 and 0.181 for respectively PSNR, SSIM and LPIPS) and our method without the distance-aware mechanism. This means that the architecture in itself (described in section 3.2), which is the sole difference between these two experiments, enables better scene reconstruction than Nerfacto.
Moreover, the addition of the distance-aware mechanism still increases the performances significantly. While this idea was initially developed as an anti-aliasing process, this experiment demonstrates that it also benefits Novel View Synthesis with little variations of distances of observation. One possible explanation comes from the experimental observation that RING-NeRF seems to better reconstruct distant areas (see figure 6 for example).
Finally, while our continuous coarse-to-fine brings slightly better results than when using the usual discrete coarse-to-fine, it still presents lower results than when not using coarse-to-fine. While the difference is not prohibitive, this observation demonstrates the presence of a trade off between quality and stability in Neural Radiance Fields. Even though using coarse-to-fine on harder setups is necessary to avoid catastrophic failure or improve consistency, as explained respectively in section 9.3 and 8.2, it is not required for simple setups such as Novel View Synthesis.
Tables 8, 9 and 10 shows per-scene results of our mono-scale experiments.
Distance-Aware Coarse-To-Fine PSNR ↑ SSIM ↑ LPIPS ↓
RING-NeRF - - 27.74 0.786 0.183
RING-NeRF - 28.11 0.802 0.155
RING-NeRF Discrete 27.84 0.788 0.168
RING-NeRF Continuous 27.98 0.793 0.168
Table 6: Novel View synthesis performances for the Mono-Scale setup on the 360 dataset.
Multi-Scale Setup.
Per-scene results can be found respectively in tables 11, 12, 13, 14, 15 and 16.
8 Few views - Section 4.3
@@ 8.1 Configuration and Evaluation Details
Regarding the model, we followed FreeNeRF’s choices as much as possible and, because there are no grid-based methods to compare ourselves to, the rest of the hyperparameters (mainly regarding the hashgrids) were derived from the configuration used in our NVS experiments. Hence, we followed the number of samples used in FreeNeRF (128 samples) and also disable the mechanisms of nerfacto dedicated to unbounded scenes : scene contraction and distortion loss. Our experiments on Nerfacto/Nerfacto+ and RING-NeRF were all done using a pyramid of 9 grids of growth factor 2, a configuration similar to our NVS experiments where we removed the last level of the highest resolution as we consider the object-centric DTU dataset simple enough to avoid to over-complexify the model. We once again use features of size 8. Based on the same observation of the dataset being simpler than the 360 dataset, we used a smaller MLP than in NVS experiments, with a hidden dimension of 64, both for density and color. However, it is important to note that because our goal was rather to demonstrate the impact of our architecture on this task against the Nerfacto+ baseline, we did not focus our research on optimizing the configuration of the models, which could prove to further increase the metrics.
Finally, we followed FreeNeRF’s density loss implementation and also use the introduced black and white prior, where we minimize the density of the $M+10$ ($5$ more samples than in FreeNeRF) first samples rather than the $M$ first when the value of the pixel is either black or white, in order to benefit from the particular form of the DTU dataset with uniform backgrounds.
@@
Table 7: Ablative Experiments on reconstruction from few viewpoints (respectively 3, 6 and 9, separated by ”/”) on the DTU dataset. The reported metrics are computed based on the mask of the object.
@@ 8.2 Ablative Experiments
First of all, the results without both of these mechanisms are, as expected, quite low. However, we notice that its results are in average better than the basic Nerfacto architecture, especially when training with 6 and 9 views. While both are low results, we see qualitatively (see figure 7) that the sole RING-NeRF architecture mostly succeeds in creating a coherent 3D geometry, although not perfect and most of all very affected by floaters artifacts. This proves that our RING-NeRF architecture is more stable than the concatenation architecture by design, even though it is still lacking on its own.
We also evaluate the model both without coarse-to-fine and without the density loss separately. Using solely the density loss surprisingly results in correct metrics. However, visualizing the predictions demonstrates that the 3D reconstruction in itself is slightly less precise than when using only coarse-to-fine, but the removal of background floaters induced by density-loss heavily increases the metrics. Moreover, the reconstruction without coarse-to-fine is less stable, with worse coherency in very hard setups, such as when observing the scene with only 3 views and with novel views very far from the training views, as illustrated in figure 8. Note that, to further demonstrate that the density loss is not the sole explanation of the results, figure 7 (b) shows the results of the nerfacto architecture coupled with the density loss, and we notice that, while a beginning of 3D consistency appears, it is still way more approximate than its RING-NeRF counterpart (e).

### v3

@@ 7.2 Result Analysis and Ablative Experiments
Nerfacto 27.09 0.779 0.181 0.45 h
PyNeRF 27.87 0.802 0.160 0.96 h
Zip-NeRF 28.06 0.808 0.154 1.10 h
Our Model 28.09 0.799 0.157 0.45 h
Multi-Scale Setup. Figure 11 illustrates how anti-aliasing models behaves with multiple distances of observations against the Nerfacto baseline. On one side, the aliased rendering of Nerfacto presents different types of artifacts at different resolutions. This is due to the NeRF’s model property to "average" the distances of observations between the trained views. This thus results in under-contrasted image at full resolution, over-contrasted image at resolution 1/4 and aliasing artifacts at resolution 1/8. On the other hand, RING-NeRF takes into account the distance in the computation of the feature and succeeds in adapting to the different resolutions.
Per-scene results can be found respectively in tables 14, 15, 16, 17, 18 and 19.
@@
8 Few Viewpoints Supervision - Section 4.3
@@ 8.1 Configuration and Evaluation Details
Regarding the model, we follow FreeNeRF’s choices as much as possible and, because there are no grid-based methods to compare ourselves to, the rest of the hyperparameters (mainly regarding the hashgrids) were derived from the configuration used in our Novel View Synthesis experiments. Hence, we followed the number of samples used in FreeNeRF (128 samples) and also disable the mechanisms of nerfacto dedicated to unbounded scenes : scene contraction and distortion loss. Our experiments on Nerfacto/Nerfacto+ and RING-NeRF are all done using a pyramid of 9 grids of growth factor 2, a configuration similar to our NVS experiments where we removed the last level of the highest resolution as we consider the object-centric DTU dataset simple enough to avoid to over-complexify the model. We once again use features of size 8. Based on the same observation of the dataset being simpler than the 360 dataset, we used a smaller MLP than in NVS experiments, with a hidden dimension of 64, both for density and color. However, it is important to note that because our goal was rather to demonstrate the impact of our architecture on this task against the Nerfacto+ baseline, we did not focus our research on optimizing the configuration of the models, which could prove to further increase the metrics.
Finally, we follow FreeNeRF’s density loss implementation and also use the introduced black and white prior, where we minimize the density of the $M+10$ ($5$ more samples than in FreeNeRF’s implementation) first samples rather than the $M$ first when the value of the pixel is either black or white, in order to benefit from the particular form of the DTU dataset with uniform backgrounds.
Table 7: Ablative Experiments on reconstruction from few viewpoints (respectively 3, 6 and 9, separated by "/") on the DTU dataset. The reported metrics are computed based on the mask of the object.
@@ 8.2 Ablative Experiments
First of all, the results without both of these mechanisms are, as expected, quite low. However, we notice that its results are in average better than the basic Nerfacto architecture, especially when training with 6 and 9 views. While both are low results, we see qualitatively (see figure 12) that the sole RING-NeRF architecture mostly succeeds in creating a coherent 3D geometry, although not perfect and most of all very affected by floaters artifacts. This proves that our RING-NeRF architecture is more stable than the concatenation architecture by design, even though it is still lacking of a way to get rid of artefacts-inducing ambiguities.
We also evaluate the model both without coarse-to-fine and without the density loss separately. Using solely the density loss surprisingly results in correct metrics. However, the visualization demonstrates that the 3D reconstruction in itself is slightly less precise than when using only coarse-to-fine, although the metrics are better as they are more impacted by the artefacts. Moreover, the reconstruction without coarse-to-fine is less stable, with worse coherency in very hard setups, such as when observing the scene with only 3 views and with novel views very far from the training views, as illustrated in figure 13. Note that, to further demonstrate that the density loss is not the sole explanation of the results, figure 12 (b) shows the results of the nerfacto architecture coupled with the density loss, and we notice that, while a beginning of 3D consistency appears, it is still way more approximate than its RING-NeRF counterpart (e).